# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ 7.3.2. Characterizations of topological pressure
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Denote $f_{X}\coloneqq f|_{X}$. Let $\phi$ be a real-valued Hölder continuous function on $X$. Then for each $x_{0}\in X\setminus\operatorname{post}(f)$, we have
@@
$P(f_{X},\phi)=$ $P(\sigma,\phi\circ\pi)$
$=$ $\lim_{n\to+\infty}\frac{1}{n}\log\sum\limits_{\omega\in\sigma^{-n}(\omega_{0})}\exp(S_{n}(\phi\circ\pi)(\omega))$
$=$ $\lim_{n\to+\infty}\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}\exp(S_{n}\phi(x)).$
The lemma is now verified. ∎
@@
By Proposition 3.11 (i), for each integer $i\in[0,n-1]$, we have $\lvert f^{i}(x)-f^{i}(y)\rvert\leq C^{\prime}d_{0}\theta^{n-i}$ since $f^{i}(x),f^{i}(y)\in f^{i}(U)\in\mathcal{U}_{n-i}$. Since $\phi$ is $\alpha$-Hölder continuous, there exists a constant $C>0$, such that $\lvert\phi(x_{1})-\phi(x_{2})\rvert\leq C\lvert x_{1}-x_{2}\rvert^{\alpha}$ for all $x_{1},\,x_{2}\in X$. So for each integer $i\in[0,n-1]$,
$\bigl{|}\phi\bigl{(}f^{i}(x)\bigr{)}-\phi\bigl{(}f^{i}(y)\bigr{)}\bigr{|}\leq C\bigl{(}C^{\prime}d_{0}\theta^{n-i}\bigr{)}^{\alpha}.$
@@
$\lvert S_{n}\phi(x)-S_{n}\phi(y)\rvert\leq\sum_{i=0}^{n-1}\bigl{|}\phi\bigl{(}f^{i}(x)\bigr{)}-\phi\bigl{(}f^{i}(y)\bigr{)}\bigr{|}\leq\sum_{i=0}^{n-1}C(C^{\prime}d_{0})^{\alpha}\cdot(\theta^{\alpha})^{n-i}\leq\frac{C(C^{\prime}d_{0}\theta)^{\alpha}}{(1-\theta^{\alpha})}.$
@@
$C_{2}^{-1}\leq\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leq C_{2},$
@@
since $f^{n}|_{U}\colon U\rightarrow U_{0}$ is an FBC map. Then by Lemma 7.15, there exists $C_{1}>0$ such that
$\exp(-C_{1})\leq\mathcal{S}_{n}(x_{0},U)/\mathcal{S}_{n}(y_{0},U)\leq\exp(C_{1}).$
Thus we get
$\exp(-C_{1})\leq\frac{\sum_{U\in\mathcal{U}_{n},f^{n}(U)=U_{0}}\mathcal{S}_{n}(x_{0},U)}{\sum_{U\in\mathcal{U}_{n},f^{n}(U)=U_{0}}\mathcal{S}_{n}(y_{0},U)}\leq\exp(C_{1}).$
@@
$C_{3}^{-1}\leq\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leq C_{3},$
@@
By Lemma 7.16, we have $C_{2}^{-1}\leq\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(u_{0})\leq C_{2}$ and $C_{2}^{-1}\leq\mathcal{S}_{n}(v_{0})/\mathcal{S}_{n}(y_{0})\leq C_{2}$ where $C_{2}>1$ is the constant from Lemma 7.16.
@@
Then by Lemma 7.13, we have $C_{0}^{-1}\leq\mathcal{S}_{n}(u_{0})/\mathcal{S}_{n}(v_{0})\leq C_{0}$. Combining the inequalities above, we have
$C_{0}^{-1}\cdot C_{2}^{-2}\leq\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leq C_{0}\cdot C_{2}^{2}.$
By putting $C_{3}\coloneqq C_{0}\cdot C_{2}^{2}$ we complete the proof of this lemma. ∎
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Denote $f_{X}\coloneqq f|_{X}$. Let $\phi$ be a real-valued Hölder continuous function on $X$. Then for each sequence $\{x_{k}\}_{k\in\mathbb{N}}$ of points in $X$,
@@
By Lemma 7.17, there exists $C_{3}>1$ such that for all $x^{\prime},x^{\prime\prime}\in X$,
$C_{3}^{-1}\leq\mathcal{S}_{n}(x^{\prime})/\mathcal{S}_{n}(x^{\prime\prime})\leq C_{3},$
where $\mathcal{S}_{n}(\cdot)$ is defined in (7.2) in Lemma 7.16. Fix an arbitrary $x_{0}\in X\setminus\operatorname{post}(f)$. Then for each $n\in\mathbb{N}$, we have
$\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}e^{S_{n}\phi(x)}-\frac{\log C_{3}}{n}\leq\frac{1}{n}\log(\mathcal{S}_{n}(x_{n}))\leq\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}e^{S_{n}\phi(x)}+\frac{\log C_{3}}{n}.$
@@
Now we are ready to prove our level-$2$ large deviation principle.
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. We denote $f_{X}\coloneqq f|_{X}$ for clarity.
@@
$\qquad=\lim_{n\to+\infty}\frac{1}{n}\biggl{(}\log\sum\limits_{y\in f^{-n}(x_{n})}e^{S_{n}(\phi+\psi)(y)}-\log\sum\limits_{y\in f^{-n}(x_{n})}e^{S_{n}\phi(y)}\biggr{)}$
@@ 7.4. Equidistribution with respect to the equilibrium state
We finish this section with an equidistribution result as a consequence of our level-$2$ large deviation principle.
First, we show a corollary of Theorem D, which gives a characterization of the measure-theoretic pressure.
Corollary 7.19.
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. We denote $f_{X}\coloneqq f|_{X}$. Let $\phi\colon X\rightarrow\mathbb{R}$ be Hölder continuous, and $\mu_{\phi}$ be the unique equilibrium state.
Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence of points in $X$. Then the following identity holds for each $\mu\in\mathcal{M}(X,f_{X})$ and each convex local basis $G_{\mu}$ of $\mathcal{P}(X)$ at $\mu$:
$h_{\mu}(f_{X})+\int\!\phi{\,\mathrm{d}}\mu=\inf\biggl{\{}\lim_{n\to+\infty}\frac{1}{n}\log\biggl{(}\sum_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\in\mathfrak{G}}\deg(f^{n};y)\cdot e^{S_{n}\phi(y)}\biggr{)}:\mathfrak{G}\in G_{\mu}\biggr{\}}.$
Proof.
Fix arbitrary $\mu\in\mathcal{M}(X,f_{X})$ and a convex local basis $G_{\mu}$ at $\mu$. By the upper semi-continuity of $h_{\mu}(f_{X})$ and the definition of $I^{\phi}(\cdot)$ in (1.1) in Theorem D, we get
(7.3) $-I^{\phi}(\mu)=\inf_{\mathfrak{G}\in G_{\mu}}\bigl{(}\sup_{\mathfrak{G}}\bigl{(}-I^{\phi}\bigr{)}\bigr{)}=\inf_{\mathfrak{G}\in G_{\mu}}\bigl{(}-\inf_{\mathfrak{G}}I^{\phi}\bigr{)}.$
Then by (1.2) in Theorem D and (7.3), we have
$-I^{\phi}(\mu)=\inf_{\mathfrak{G}\in G_{\mu}}\Biggl{\{}\lim_{n\to+\infty}\frac{1}{n}\log\frac{\sum\limits_{y\in f^{-n}(x_{n}),\,W_{n}(y)\in\mathfrak{G}}\deg(f^{n};y)\cdot e^{S_{n}\phi(y)}}{Z_{n}(\phi)}\Biggr{\}},$
where $Z_{n}(\phi)\coloneqq\sum_{y\in f^{-n}(x_{n})}\deg(f^{n};y)\exp(S_{n}\phi(y))$. Then since $\lim_{n\to+\infty}\frac{1}{n}Z_{n}(\phi)=P(f_{X},\phi)$ (see Lemma 7.18), the proof is complete. ∎
We now prove the equidistribution result.
Proof of Theorem E.
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. We denote $f_{X}\coloneqq f|_{X}$ for clarity. Let $\phi\colon X\rightarrow\mathbb{R}$ be Hölder continuous, and $\mu_{\phi}$ be the unique equilibrium state.
We denote
$Z^{+}_{n}(\mathfrak{G})$ $\coloneqq\sum_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\in\mathfrak{G}}\deg(f^{n};y)\cdot\exp(S_{n}\phi(y)),$
$Z^{-}_{n}(\mathfrak{G})$ $\coloneqq\sum_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\notin\mathfrak{G}}\deg(f^{n};y)\cdot\exp(S_{n}\phi(y))$
for each $n\in\mathbb{N}$ and each open set $\mathfrak{G}\subseteq\mathcal{P}(X)$.
Let $G_{\mu_{\phi}}$ be a convex local basis of $\mathcal{P}(X)$ at $\mu_{\phi}$. Fix an arbitrary convex open set $\mathfrak{G}\in G_{\mu_{\phi}}$.
By the uniqueness of the equilibrium state in Proposition 7.8, for each $\mu\in\mathcal{P}(X)\setminus\{\mu_{\phi}\}$, we get that
$h_{\mu}(f_{X})+\int\phi{\,\mathrm{d}}\mu<P(f_{X},\phi).$
Then by Corollary 7.19, there exists an open neighborhood $\mathfrak{U}_{\mu}\subseteq\mathcal{P}(X)\setminus\{\mu_{\phi}\}$ of $\mu$ such that
$\lim_{n\to+\infty}\frac{1}{n}Z_{n}^{+}(\mathfrak{U}_{\mu})<P(f_{X},\phi).$
Thus there exist numbers $a_{\mu}<P(f_{X},\phi)$ and $N_{\mu}\in\mathbb{N}$, such that for each $n>N_{\mu}$,
$Z^{+}_{n}(\mathfrak{U}_{\mu})\leq\exp(na_{\mu}).$
Since $\mathcal{P}(X)$ is $*{}^{*}$ compact, so is $\mathcal{P}(X)\setminus\mathfrak{G}$. Thus there exists a finite set $\{\mu_{i}:i\in I\}\subseteq\mathcal{P}(X)\setminus\mathfrak{G}$, where $I$ is a finite index set, such that
$\mathcal{P}(X)\setminus\mathfrak{G}\subseteq\bigcup_{i\in I}\mathfrak{U}_{\mu_{i}}.$
Put $a\coloneqq\max\{a_{\mu_{i}}:i\in I\}$. Note that $a<P(f_{X},\phi)$. Then we have
(7.4) $Z^{-}_{n}(\mathfrak{G})\leq\sum_{i\in I}Z^{+}_{n}(\mathfrak{U}_{\mu_{i}})\leq\operatorname{card}(I)\exp(na)$
for each $n>\max\{N_{\mu_{i}}:i\in I\}$.
Applying Corollary 7.19 with $\mu=\mu_{\phi}$, we get that
$P(f_{X},\phi)\leq\lim_{n\to+\infty}\frac{1}{n}\log\bigl{(}Z^{+}_{n}(\mathfrak{G})\bigr{)}.$
On the other hand, by Lemma 7.18, we have
$P(f_{X},\phi)=\lim_{n\to+\infty}\frac{1}{n}\log\bigl{(}Z_{n}(\phi)\bigr{)}\geq\lim_{n\to+\infty}\frac{1}{n}\log\bigl{(}Z^{+}_{n}(\mathfrak{G})\bigr{)},$
so we get that
$P(f_{X},\phi)=\lim_{n\to+\infty}\frac{1}{n}\log\bigl{(}Z^{+}_{n}(\mathfrak{G})\bigr{)}.$
Choose numbers $b\in(a,P(f_{X},\phi))$ and $N\geq\max\{N_{\mu_{i}}:i\in I\}$ such that for each integer $n>N$,
(7.5) $Z^{+}_{n}(\mathfrak{G})\geq\exp(nb).$
Combining (7.4) with (7.5), we get
(7.6) $\lim_{n\to+\infty}\frac{Z^{-}_{n}(\mathfrak{G})}{Z^{+}_{n}(\mathfrak{G})}\leq\lim_{n\to+\infty}\frac{\operatorname{card}(I)\exp(na)}{\exp(nb)}=0.$
Put
$\nu^{\prime}_{n}\coloneqq\sum\limits_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\in\mathfrak{G}}\frac{\deg(f^{n};y)\cdot\exp(S_{n}\phi(y))}{Z^{+}_{n}(\mathfrak{G})}W_{n}(y),$
where $W_{n}(y)\coloneqq\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}(y)}$. Then for each $n\in\mathbb{N}$, we have
$\nu_{n}=\frac{Z^{+}_{n}(\mathfrak{G})}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}\nu^{\prime}_{n}+\sum\limits_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\notin\mathfrak{G}}\frac{\deg(f^{n};y)\cdot\exp(S_{n}\phi(y))}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}W_{n}(y).$
By (7.6) we know that $\lim_{n\to+\infty}\frac{Z^{+}_{n}(\mathfrak{G})}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}=1$ and that the total variation
$\biggl{\|}\sum\limits_{y\in f^{-n}(x_{n}),\atop W_{n}(y)\notin\mathfrak{G}}\frac{\deg(f^{n};y)\cdot\exp(S_{n}\phi(y))}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}W_{n}(y)\biggr{\|}$
$\qquad\leq\frac{\sum\limits_{y\in f^{-n}(x_{n}),\,W_{n}(y)\notin\mathfrak{G}}\deg(f^{n};y)\cdot\exp(S_{n}\phi(y))\|W_{n}(y)\|}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}$
$\qquad\leq\frac{Z^{-}_{n}(\mathfrak{G})}{Z^{+}_{n}(\mathfrak{G})+Z^{-}_{n}(\mathfrak{G})}\longrightarrow 0$
as $n$ tends to 00. Thus a measure is a subsequential limit of $\{\nu_{n}\}_{n\in\mathbb{N}}$ if and only if it is a subsequential limit of $\{\nu^{\prime}_{n}\}_{n\in\mathbb{N}}$. Note that for each $n\in\mathbb{N}$, $\nu^{\prime}_{n}$ is a convex combination of measures in the convex set $\mathfrak{G}$, so $\nu^{\prime}_{n}\in\mathfrak{G}$. Hence each subsequential limit of $\{\nu_{n}\}_{n\in\mathbb{N}}$ lies in $\overline{\mathfrak{G}}$.
Since $\mathfrak{G}\in G_{\mu_{\phi}}$ is arbitrary, we get that each subsequential limit of $\{\nu_{n}\}_{n\in\mathbb{N}}$ is $\mu_{\phi}$. Therefore $\nu_{n}$ converges to $\mu_{\phi}$ in the $*{}^{*}$ topology as $n$ tends to infinity. ∎

### v3

@@ 7.3.2. Characterizations of topological pressure
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\phi$ be a real-valued Hölder continuous function on $X$. Then for each $x_{0}\in X\setminus\operatorname{post}(f)$, we have
@@
$P(f_{X},\phi)=P(\sigma,\phi\circ\pi)=$ $\lim_{n\to+\infty}\frac{1}{n}\log\sum\limits_{\omega\in\sigma^{-n}(\omega_{0})}\exp(S_{n}(\phi\circ\pi)(\omega))$
$=$ $\lim_{n\to+\infty}\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}\exp(S_{n}\phi(x)).\qed$
@@
By Proposition 3.11 (i), for each integer $i\in[0,n-1]$, we have $\lvert f^{i}(x)-f^{i}(y)\rvert\leqslant C^{\prime}d_{0}\theta^{n-i}$ since $f^{i}(x),f^{i}(y)\in f^{i}(U)\in\mathcal{U}_{n-i}$. Since $\phi$ is $\alpha$-Hölder continuous, there exists a constant $C>0$, such that $\lvert\phi(x_{1})-\phi(x_{2})\rvert\leqslant C\lvert x_{1}-x_{2}\rvert^{\alpha}$ for all $x_{1},\,x_{2}\in X$. So for each integer $i\in[0,n-1]$,
$\bigl{\lvert}\phi\bigl{(}f^{i}(x)\bigr{)}-\phi\bigl{(}f^{i}(y)\bigr{)}\bigr{\rvert}\leqslant C\bigl{(}C^{\prime}d_{0}\theta^{n-i}\bigr{)}^{\alpha}.$
@@
$\lvert S_{n}\phi(x)-S_{n}\phi(y)\rvert\leqslant\sum_{i=0}^{n-1}\bigl{\lvert}\phi\bigl{(}f^{i}(x)\bigr{)}-\phi\bigl{(}f^{i}(y)\bigr{)}\bigr{\rvert}\leqslant\sum_{i=0}^{n-1}C(C^{\prime}d_{0})^{\alpha}\cdot(\theta^{\alpha})^{n-i}\leqslant\frac{C(C^{\prime}d_{0}\theta)^{\alpha}}{(1-\theta^{\alpha})}.$
@@
$C_{2}^{-1}\leqslant\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leqslant C_{2},$
@@
since $f^{n}|_{U}\colon U\rightarrow U_{0}$ is an FBC map. Then it is easy to see from Lemma 7.15 that
$\exp(-C_{1})\leqslant\mathcal{S}_{n}(x_{0},U)/\mathcal{S}_{n}(y_{0},U)\leqslant\exp(C_{1}),$
where $C_{1}>0$ is the constant from Lemma 7.15 Thus we get
$\exp(-C_{1})\leqslant\frac{\sum_{U\in\mathcal{U}_{n},f^{n}(U)=U_{0}}\mathcal{S}_{n}(x_{0},U)}{\sum_{U\in\mathcal{U}_{n},f^{n}(U)=U_{0}}\mathcal{S}_{n}(y_{0},U)}\leqslant\exp(C_{1}).$
@@
$C_{3}^{-1}\leqslant\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leqslant C_{3},$
@@
By Lemma 7.16, we have $C_{2}^{-1}\leqslant\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(u_{0})\leqslant C_{2}$ and $C_{2}^{-1}\leqslant\mathcal{S}_{n}(v_{0})/\mathcal{S}_{n}(y_{0})\leqslant C_{2}$ where $C_{2}>1$ is the constant from Lemma 7.16.
@@
Then, by Lemma 7.13, we have $C_{0}^{-1}\leqslant\mathcal{S}_{n}(u_{0})/\mathcal{S}_{n}(v_{0})\leqslant C_{0}$. Combining the inequalities above, we have
$C_{0}^{-1}\cdot C_{2}^{-2}\leqslant\mathcal{S}_{n}(x_{0})/\mathcal{S}_{n}(y_{0})\leqslant C_{0}\cdot C_{2}^{2}.$
By putting $C_{3}\coloneqq C_{0}\cdot C_{2}^{2}$, we complete the proof of this lemma. ∎
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\phi$ be a real-valued Hölder continuous function on $X$. Then for each sequence $\{x_{k}\}_{k\in\mathbb{N}}$ of points in $X$,
@@
By Lemma 7.17,
$C_{3}^{-1}\leqslant\mathcal{S}_{n}(x^{\prime})/\mathcal{S}_{n}(x^{\prime\prime})\leqslant C_{3}$
holds for all $x^{\prime},x^{\prime\prime}\in X$, where $C_{3}>1$ is the constant from Lemma 7.17, and $\mathcal{S}_{n}(\cdot)$ is defined in (7.2) in Lemma 7.16. Fix an arbitrary $x_{0}\in X\setminus\operatorname{post}(f)$. Then for each $n\in\mathbb{N}$, we have
$\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}e^{S_{n}\phi(x)}-\frac{\log C_{3}}{n}\leqslant\frac{1}{n}\log(\mathcal{S}_{n}(x_{n}))\leqslant\frac{1}{n}\log\sum\limits_{x\in f^{-n}(x_{0})}e^{S_{n}\phi(x)}+\frac{\log C_{3}}{n}.$
@@
Now, we are ready to prove our level-$2$ large deviation principle.
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2.
@@
$\qquad=\lim_{n\to+\infty}\frac{1}{n}\biggl{(}\log\sum\limits_{y\in f^{-n}(x_{n})}\deg(f^{n};y)\cdot e^{S_{n}(\phi+\psi)(y)}-\log\sum\limits_{y\in f^{-n}(x_{n})}\deg(f^{n};y)\cdot e^{S_{n}\phi(y)}\biggr{)}$
@@
We finish this paper with an equidistribution result as a consequence of our level-$2$ large deviation principle. The proof follows from similar arguments in the proof of [Li15, Corollary 1.7].
Corollary 7.19 (Equidistribution).
  1. (1)
    $\mathfrak{X}_{0}$ and $\mathfrak{X}_{1}$ are strongly path-connected.
  2. (2)
    The branch set $B_{f}$ is finite.
Let $\phi$ be a real-valued Hölder continuous function on $X$, and $\mu_{\phi}$ be the unique equilibrium state for $f|_{X}$ and potential $\phi$. Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence of points in $X$. Consider the following sequence of Borel probability measures on $X$:
where $S_{n}\phi$ is defined in (2.1). Then $\nu_{n}$ converges to $\mu_{\phi}$ in the weak topology as $n\to+\infty$.